OpenStudy (anonymous):

Write the equation of a line with a slope of −2 and a y-intercept of 5. −2x + y = 5 2x − y = 5 y = −2x + 5 y = 5x − 2

OpenStudy (anonymous):

helpp

imqwerty (imqwerty):

y=-2x+5 is the answer

OpenStudy (anonymous):

can u explain that lol

imqwerty (imqwerty):

we just have to put the values of m and c i.e, slope and y intercept of the line respectively in the slope intercept form i.e, y=mx+c

OpenStudy (anonymous):

oo ok

OpenStudy (anonymous):

can u help me with like 2 more problems

imqwerty (imqwerty):

ok

OpenStudy (anonymous):

Shameka is tossing acorns into a hole. She started with 800 acorns in the hole. On average, she gets 75 acorns per hour in the hole. Write an equation that represents the amount of acorns, A, in the hole, where h is the number of hours. A = 75h – 800 A = 75h + 800 A = −800h – 75 A = −800h + 75

OpenStudy (haseeb96):

y=mx + c this is the formula here m is the slope and c is the y intercept. as we know m=-2 and c=5 now just put these values in the above formula and get the answer

OpenStudy (anonymous):

Shameka is tossing acorns into a hole. She started with 800 acorns in the hole. On average, she gets 75 acorns per hour in the hole. Write an equation that represents the amount of acorns, A, in the hole, where h is the number of hours. A = 75h – 800 A = 75h + 800 A = −800h – 75 A = −800h + 75

imqwerty (imqwerty):

the answer is 75h+800 the total number of acorns in hole = number of acorns already in hole(800) + number of acorns added by her till that time(75h)

OpenStudy (anonymous):

Write the equation of the line that passes through (1, 5) and (–2, 14) in slope-intercept form. y = 3x + 2 y = 3x + 8 y = –3x – 2 y = –3x + 8

OpenStudy (anonymous):

can someone help meeee